OpenStudy (anonymous):

whats square root 20v^6??

OpenStudy (tkhunny):

\(\sqrt{20v^{6}}\) Find the perfect squares \(\sqrt{5\cdot 4\cdot v^{6}}\) Attack them one at a time. \(\sqrt{5}\sqrt{4}\sqrt{v^{6}} = \sqrt{5}\cdot 2\cdot \left|v^{3}\right|\)

OpenStudy (anonymous):

So the answer is \[\sqrt{5}\]*2*|v^3| ?

OpenStudy (anonymous):

\[\sqrt{5}\times2timesv ^{3}\]

OpenStudy (tkhunny):

Absolute Value of v^3.

OpenStudy (anonymous):

wait so whats the answer? hah sorry D:

OpenStudy (tkhunny):

I already wrote it up above. Now, let's see you do one. Only one for free.

OpenStudy (anonymous):

But your answer doesn't work...

OpenStudy (tkhunny):

The answer is correct as I have stated it. However, someone may have overlooked the absolute values. Try it without them. It will be wrong, but it may take it. Next time, let's try YOUR answer and see if it works.

OpenStudy (anonymous):

it says assuming v represents a positive real number

OpenStudy (anonymous):

i got √5√4 but didnt know what to do with v^6

OpenStudy (tkhunny):

Very good. That is wonderful information. You did not provide it, before. This makes all the difference. Now, we have \(2v^{3}\sqrt{5}\). Please provide the ENTIRE problem statement and please provide your best work.

OpenStudy (tkhunny):

If v > 0, we have \(\sqrt{v^{6}} = [v^{6}]^{1/2} = v^{6\cdot (1/2)} = v^{3}\)

OpenStudy (anonymous):

Okay since v^6 reduces to v^3, then its just 2v^3√5

OpenStudy (tkhunny):

You tell me. You're the one who has the software that thinks it knows stuff. Did we get what it wants, or not? It is very important to note that "v^{6} reduces to v^3" ONLY because we know v > 0.

OpenStudy (anonymous):

yes! that one worked. thank you!!